OpenStudy (anonymous):

Use induction to show that 1+3+5+...+(2N-1)=N^2

OpenStudy (anonymous):

Base case \(n=1\). There is only one term to add so... We need to show that \(2(\color{red}{1})-1=(\color{red}1)^2\)

OpenStudy (misty1212):

hi? do you know how to do a proof by induction? summation ones are the easiest

OpenStudy (anonymous):

so, it would be 1+3+5+..+(2(1)-1)-1=1^2 @wio

OpenStudy (anonymous):

You assume:\[ 1+3+5+\ldots+(2N-1)=N^2 \]And then add the next term:\[ 1+3+5+\ldots+(2N-1)+(2N+1)=N^2 + (2N+1) \]And show that this somehow becomes: \[ 1+3+5+\ldots+(2(N+1)-1)=(N+1)^2 \]

OpenStudy (anonymous):

@sar12389 Yeah, but you don't have the \(1+3+5\) because in the base case, there is only the \(1\) term because \(n\) is too small to reach those terms

OpenStudy (anonymous):

so it would just be (2(n+1)-1)=n+1^2 @wio

OpenStudy (anonymous):

@wio ?^